OpenStudy (anonymous):

it is unknown that y is proportional to the square of x. Find by what factor y is multiplied if x is doubled...

OpenStudy (anonymous):

it is unknown, or it is known?

OpenStudy (anonymous):

unknown

OpenStudy (anonymous):

OK, then the question is impossible.

OpenStudy (anonymous):

are you sure about that?... the answer is 1/2 ; 2 .. it says that at the back of the textbook...

OpenStudy (anonymous):

underneath the question, it says if y=18 when x=3, find ghe value of y when x=10

OpenStudy (anonymous):

if the question is "it is known that y is proportional to the square of x," then if x doubles, y is multiplied by \[\sqrt{2}\]if the question is "it is unknown that y is proportional to the square of x" then that makes no sense whatsoever.

OpenStudy (anonymous):

how about this one: given that y varies as the cube of x, Find the factor by which y is multioplied if x is doubled... underneath the question it says : by what factor is y multiplied if x is multiplied by 1/3

OpenStudy (anonymous):

if x is doubled, y goes up by \[\sqrt[3]{2}\]if x is multiplied by 1/3, y is multiplied by \[\sqrt[3]{\frac{1}{3}}\]

OpenStudy (anonymous):

how did you figure that out?...can you explain your working out??

OpenStudy (anonymous):

y varies as \[x^3\]so x varies as \[\sqrt[3]{y}\]

OpenStudy (anonymous):

y varies as x cubed just means \[y=ax^3\]where a is a constant that you don't know. you can solve for x and get\[x=\frac{1}{\sqrt[3]{a}}\sqrt[3]{y}\]where the a part is still a constant that you don't care about. So as x gets multiplied by a factor, y gets multiplied by the cube root of that factor.

OpenStudy (anonymous):

aah.... thank you soooo much for your help! :)